A string clamped at both ends, vibrates in a pattern of six loops at a frequency of 240 Hz. What frequency will cause it to vibrate in four loops?

1
Expert's answer
2021-12-20T10:29:30-0500

Explanations & Calculations


  • Hello Ahmad, explanation to this question also refers to that made for the question Q279421.
  • For a string clamped at both ends, the formula "\small f = \large\frac{nv}{2L}" gives the frequency of each mode of vibration.
  • Hoping you are conversant with the derivation of that formula, it could be said that, that "\small n" represents the number of loops present while in a vibration.
  • Therefore, for the 6 loops,

"\qquad\qquad
\begin{aligned}
\small 240&=\small \frac{6v}{2L}\implies 80=\frac{v}{L}\cdots(1)
\end{aligned}"

  • For the frequency at 4 loops,

"\qquad\qquad
\begin{aligned}
\small f&=\small \frac{4v}{2L}\implies f=2\Big(\frac{v}{L}\Big)
\end{aligned}"

  • Now you know the value of the "\small v/L" ratio, you can give it a try to get the answer.